Simplify the expressions and solve the equations

Student SIdult 10. Věc+6 = v3c-9 S.a.p. 11. (5 +687) Sap. LE3
Answer #1

₃6 (+6 = 136-9 on Squaring both sides. - +6 = 3C-9 > 6+9=3C- - 50 = 15 : [C=6 ANS. til (5+6+5) = 5+677)*+2x5*637 = 25+252 + b